OpenStudy (anonymous):

Express 3^2 = x as a logarithmic equation.

OpenStudy (nirmalnema):

3^2=x x=9 log of 9 to the base 3 = 2

OpenStudy (anonymous):

What is the solution of log3x+4 125 = 3

OpenStudy (nirmalnema):

what is b/w 4 and 125

OpenStudy (anonymous):

what

OpenStudy (anonymous):

log(3x – 2) 125 = 3 log(3x - 2) = (3/125) 10^(log(3x-2)) = 10^(3/125) = 1.116863 3x - 2 = 1.116863 3x = 3.116863 x = 1.038954